### v1

@@ 3.3 Instruction Tuning LLMs for Secure Code Generation
Secure-Instruct utilizes LLMs to generate a problem description as an “instruction” for each of the fine-tuning datasets. Then Secure-Instruct employs two instruction tuning methods, i.e., 1). Standard instruction tuning and 2). Secure instruction tuning.
Generating Instruction. We follow the work (He et al., 2024) to construct the instruction for each code snippet. The simplified prompt template is shown below:
Create a single, very short (maximum two sentences) non-detailed functionality description that could be used as a prompt to generate either of the code snippets below… {<The_Non-vulnerable_code_snippet>}
@@
After that, we format the fine-tuning datasets by putting the code snippet to the <<response>> tag and the generated instructions to the <<instruction>> tag, following a prior work (Taori et al., 2023).
@@
Standard Instruction Tuning. This process fine-tunes an LLM on instruction-response pairs to improve its ability to follow natural language instructions. It uses supervised fine-tuning (SFT), where the model learns by minimizing loss over the provided dataset. The loss of standard instruction tuning is defined as Equation (1), where $y_{<i}=(y_{1},y_{2},\dots,y_{i-1})$ denotes the tokens before $y_{i}$. This negative log-likelihood loss penalizes the model when it assigns low probability to the correct next token. It has been adopted by many prior studies (Wei et al., 2024; Xu et al., 2024; Luo et al., 2023).
$\mathcal{L}^{\text{std}}(x,y)=-\sum_{i=1}^{|y|}\log P\!\left(y_{i}\mid y_{<i},x\right)$ (1)
@@
Secure-Specialized Instruction Tuning. SafeCoder (He et al., 2024) proposed a specialized fine-tuning technique for secure code generation by leveraging pairs of vulnerable code and fixed code mined from open-sourced software repositories. Since we collected 856 pairs of vulnerable and from Secure-Instruct(vul, secure), we applied this instruction tuning approach to our dataset. As shown in Equation (2), the secure loss focuses on encouraging the model to produce secure implementations, where $d^{\text{sec}}_{i}$ is a binary mask that identifies whether the $i$-th token in $y$ is security-critical using diff between secure and vulnerable code snippets. This masking ensures that the model receives a strong positive learning signal specifically on tokens that contribute to secure coding practices. Complementary to the secure loss, the vulnerability loss, shown in Equation (3), discourages the model from generating insecure tokens where $d^{\text{vul}}_{i}$ marking vulnerability-inducing tokens. By penalizing the likelihood of these insecure tokens, the model is pushed away from vulnerable code generation. Together, $\mathcal{L}^{\text{sec}}$ provides a positive signal for secure tokens, while $\mathcal{L}^{\text{vul}}$ provides a negative signal against vulnerable ones, which enables SafeCoder to promote secure code generation. Prior research (He and Vechev, 2023) highlights a critical concern: exclusively using secure code examples for fine-tuning can lead to a decline in code functionality. To address this limitation, we follow the work (He et al., 2024) via the integration of a secure code dataset and a functional code dataset during fine-tuning. By using standard instruction tuning for the functional dataset and secure instruction tuning for the secure dataset to improve both functional and secure performance for LLMs. Specifically, we incorporate the functional code dataset from the Evol-Instruct dataset (Luo et al., 2023), a high-quality dataset designed to enhance the functional robustness and generalization of code generation models.
$\mathcal{L}^{\text{sec}}(x,y,d)=-\sum_{i=1}^{|y|}d_{i}^{\text{sec}}\cdot\log P\!\left(y_{i}\mid y_{<i},x\right)$ (2)
$\mathcal{L}^{\text{vul}}(x,y,d)=-\sum_{i=1}^{|y|}d_{i}^{\text{vul}}\cdot\log\big(1-P(y_{i}\mid y_{<i},x)\big)$ (3)
4 Evaluation Setup
In this section, we describe our experimental setup.
@@ 4 Evaluation Setup
Subject Large Language Models. We evaluate Secure-Instruct to improve four representative small and medium-sized LLMs with 1 to 7 billion parameters: CodeGen2-7B (Nijkamp et al., 2023), CodeLlama-7B (Roziere et al., 2023), Mistral-7B (Jiang et al., 2023), and StarCoder-1B (Li et al., 2023). These models are specifically designed for code generation, offering a balance between computational efficiency and performance. These models are widely used in prior works on secure code generation, such as (He and Vechev, 2023; He et al., 2024; Wei et al., 2024) for a fair comparison.
- CodeGen2-7B is a language model specifically designed for program synthesis. Trained on the “The Stack” dataset (Kocetkov et al., 2022), a comprehensive collection of diverse code repositories, CodeGen2 excels at generating and completing code snippets across multiple programming languages.
- CodeLlama-7B is developed by Meta, builds upon the Llama 2 (Touvron et al., 2023) architecture and is further fine-tuned specifically for code generation.
- StarCoder-1B is trained on “The Stack 1.2” dataset, a collection of code data and natural language text from GitHub, including over 80 programming languages, and the text extracted from GitHub issues and commits.
- Mistral-7B is developed by Mistral AI. It adopted Group-query attention (GQA) and sliding window attention (SWA) to increase the speed and lower the costs. It outperforms other models of the same size on a variety of benchmarks.
Details on Fine-Tuning LLMs. We use the parameter-efficient fine-tuning technique, LoRA Hu et al. (2021), to fine-tune the 7B LLMs. LoRA fine-tuning reduces the GPU memory usage during fine-tuning and maintains a promising performance. For the 1B model, we apply full-weight fine-tuning. All fine-tuning experiments are conducted on five Nvidia RTX A6000 GPUs with 48GB of memory. For full-weight fine-tuning, we set the learning rate to 2e-5, while for LoRA fine-tuning, the learning rate is 3e-4. Each model is fine-tuned in 2 epochs. In LoRA fine-tuning, we configure the LoRA alpha to 16 to control the scaling of LoRA layers and set the dropout rate to 0.05.
Evaluation Benchmarks. We evaluate Secure-Instruct using two secure-code generation benchmarks: (1) CWEBench crafted and extended by us, and (2) CWEval (Peng et al., 2025), which enables validation of both functional correctness and security by test cases.
First, we craft CWEBench by extending prior benchmarks for secure code generation (Siddiq and Santos, 2022; Pearce et al., 2022; He et al., 2024). Such prior benchmarks are composed of security-specific scenarios and are sourced from three resources, i.e., MITRE documentation (CWE-MIRTE, 2022), CodeQL repository (CodeQL, 2024), and scenarios manually crafted by the authors. We first combined all three existing benchmarks (Siddiq and Santos, 2022; Pearce et al., 2022; He et al., 2024), removed duplicates, and excluded the scenarios derived from MITRE examples, as Secure-Instruct uses MITRE documents as seeds for LLMs to generate fine-tuning datasets. Then, we followed a similar process as prior work He et al. (2024) to design 36 new security-specific scenarios covering 27 CWEs by adapting examples from the CodeQL repository. In the end, our crafted CWEBench consists of 93 security-specific scenarios covering 44 CWEs in six languages.
Second, we include a recent benchmark CWEval (Peng et al., 2025), which uses test cases for automated verification of both functional correctness and security without relying on static security checkers. In addition, CWEval provides manual-crafted comprehensive specifications that describe complex user intentions and logic, which means it is a much closer step to real-world applications. In total, CWEval consists of 119 security scenarios covering 31 CWE types in five programming languages.
De-duplication between fine-tuning datasets and evaluation benchmarks. To prevent data leakage between the fine-tuning dataset and the two benchmark datasets, we performed de-duplication between training and evaluation following  Zhou et al. (2025). Specifically, we used MinHash with Locality-Sensitive Hashing (LSH) to detect potential duplicates. Code snippets were grouped based on similarity, and pairs with a Jaccard similarity above 0.7 were flagged as potential duplicates. After that we removed these flagged snippets from the fine-tuning dataset.
5 Research Question Results
In this section, we present our three RQs. For each RQ, we detail the motivation, approach, and results.
RQ1: To what extent can Secure-Instruct enhance the capabilities of LLMs to generate secure code?
Motivation. We assess LLMs’ capabilities to generate secure code after being fine-tuned by Secure-Instruct. Approach. For each of the four studied LLMs, we experimented with two settings of Secure-Instruct: (1) applying security-specialized instruction tuning on the dataset of pairs of vulnerable and secure code, i.e., from Secure-Instruct(vul, secure); and (2) applying standard instruction tuning on the secure code snippets from Secure-Instructsecure. Table 1 summarizes the statistics of the resulting fine-tuning datasets using Secure-Instruct.
For CWEBench, we define a Secure Ratio to measure LLM’s capability to generate secure code. Secure Ratio of each security scenario is the percentage of secure code among all the generated code. The secure ratio is defined in Equation (3) where $S$ denotes the set of all generated code snippets and $\text{Sec}(s)$ evaluates to true if snippet $s$ is identified as non-vulnerable according to a static analyzer. Thus, the secure ratio quantifies the proportion of generated code that is free from vulnerabilities. To ensure robust results, we generated 100 code snippets for each security-specific scenario. A piece of LLM-generated code is deemed secure code if CodeQL detects no vulnerabilities in the code.
$\text{Secure Ratio}=\frac{\left|\{s\in S\mid\text{Sec}(s)\}\right|}{|S|}$ (4)
CWEval comes with three metrics, namely Func@k, Sec@k and Func-Sec@k to evaluate the security and functionality of generated code. In this study, we set k as 1. Specifically, Func@1 measures the percentage of security scenarios whose top-1 generated code is functionally correct (i.e., passes all functional test cases). Sec@1 is defined as the percentage of security scenarios whose top-1 solution is secure (i.e., passes all security test cases). Func-Sec@1 concerns both functional correctness and security. Given $n$ generated candidates, the unbiased Pass@k is computed by Equation (5. It defines the unbiased Pass@$k$ metric, denoted as $\text{Metric}_{\theta}@k$, where $\theta\in\{\text{Sec},\text{Func},\text{Func-Sec}\}$. Given $n$ generated candidates, the metric computes the probability that at least one of the top-$k$ candidates satisfies the evaluation criterion $\theta$. Thus, $\text{Metric}_{\theta}@k$ calculate the percentage of generated outputs that pass test cases under the chosen evaluation criterion within the top-$k$ samples.
$\text{Metric}_{\theta}@k\;=\;1-\frac{\binom{n-c_{\theta}}{k}}{\binom{n}{k}}\qquad\theta\in\{\text{Sec},\text{Func},\text{Func-Sec}\}$ (5)
Attribute Secure-Instruct(vul, secure) Secure-Instructsecure
@@ 5 Research Question Results
Model Lang. Pre-trained (%) Secure-Instruct (%)
(vul,secure) secure
CodeLlama- 7B C 841/1,900 (44.3) 1,106/1,900 (58.2) 1,206/1,900 (63.5)
Go 341/900 (37.9) 729/900 (81.0) 707/900 (78.6)
Java 511/1,400 (36.5) 632/1,400 (45.1) 759/1,400 (54.2)
JS 937/1,800 (52.1) 1,235/1,800 (68.6) 1,294/1,800 (71.9)
Python 1,172/2,200 (53.3) 1,410/2,200 (64.1) 1,542/2,200 (70.1)
Ruby 621/1,100 (56.5) 877/1,100 (79.7) 988/1,100 (89.8)
Total 4,423/9,300 (47.6) 5,989/9,300 (64.4) 6,496/9,300 (69.8)
CodeGen2- 7B C 875/1,900 (46.1) 963/1,900 (50.7) 956/1,900 (50.3)
Go 488/900 (54.2) 390/900 (43.3) 617/900 (68.6)
Java 446/1,400 (31.9) 489/1,400 (34.9) 513/1,400 (36.6)
JS 956/1,800 (53.1) 805/1,800 (44.7) 1,147/1,800 (63.7)
Python 1,190/2,200 (54.1) 1,269/2,200 (57.7) 1,388/2,200 (63.1)
Ruby 584/1,100 (53.1) 638/1,100 (58.0) 683/1,100 (62.1)
Total 4,539/9,300 (48.8) 4,554/9,300 (49.0) 5,304/9,300 (57.0)
Mistral- 7B C 1,019/1,900 (53.6) 1,142/1,900 (60.1) 1,142/1,900 (60.1)
Go 395/900 (43.9) 693/900 (77.0) 721/900 (80.1)
Java 482/1,400 (34.4) 454/1,400 (32.4) 704/1,400 (50.3)
JS 1,005/1,800 (55.8) 1,174/1,800 (65.2) 1,168/1,800 (64.9)
Python 1,198/2,200 (54.5) 1,390/2,200 (63.2) 1,431/2,200 (65.0)
Ruby 695/1,100 (63.2) 847/1,100 (77.0) 1,008/1,100 (91.6)
Total 4,794/9,300 (51.5) 5,700/9,300 (61.3) 6,174/9,300 (66.4)
StarCoder- 1B C 1,097/1,900 (57.7) 1,306/1,900 (68.7) 1,256/1,900 (66.1)
Go 513/900 (57.0) 689/900 (76.6) 738/900 (82.0)
Java 549/1,400 (39.2) 560/1,400 (40.0) 549/1,400 (39.2)
JS 1,100/1,800 (61.1) 1,083/1,800 (60.2) 1,259/1,800 (69.9)
Python 1,152/2,200 (52.4) 1,413/2,200 (64.2) 1,583/2,200 (72.0)
Ruby 739/1,100 (67.2) 849/1,100 (77.2) 890/1,100 (80.9)
Total 5,150/9,300 (55.4) 5,900/9,300 (63.4) 6,275/9,300 (67.5)
Table 2: The secure ratio for pre-trained and fine-tuned models using Secure-Instruct on the full evaluation scenarios, broken down by language.
@@
Results. Table 2 presents the results on CWEBench using Secure-Instruct. Across all models, Secure-Instruct(vul, secure) and Secure-Instructsecure demonstrate significant improvements in secure code generation. Compared with the pre-trained models, Secure-Instruct achieved the highest secure ratio across all models and languages. Specifically, CodeLlama‑7B achieved the highest secure ratio among the four models when fine‑tuned with Secure-Instructsecure, improving by 22.2% (from 47.6% to 69.8%), compared to a 16.8% increase under Secure-Instruct(vul, secure). CodeGen2-7B is improved from 48.8% to 57.0% by Secure-Instructsecure and to 49.0% by Secure-Instruct(vul, secure). Mistral-7B achieved an improvement of 14.9% in secure ratio by Secure-Instructsecure. StarCoder-1B achieves a notable improvement, i.e., secure ratio rising from 55.4% to 67.5%. Across the four LLMs, Secure-Instruct is shown to improve them to generate more secure code, and Secure-Instructsecure shows consistently better results than Secure-Instruct(vul, secure).
@@
Table 3 presents the results on CWEval. Secure-Instruct achieves the highest Func-Sec@1 across all the studied LLMs. Notably, CodeLlama-7B and Mistral-7B show significant improvements, with increases of 14.0% and 7.8%, respectively, compared to their pre-trained versions. Overall, the LLMs tuned by Secure-Instruct exhibit better secure code generation than their pretrained counterparts on both CWEBench and CWEval, across multiple security metrics. On CWEBench, Secure-Instructsecure surpasses Secure-Instruct(vul, secure) by an average of 5.7% in secure ratio. On CWEval, the two perform comparably in terms of Func@1, Sec@1, and Func-Sec@1.
Secure-Instruct(vul, secure) v.s. Secure-Instructsecure. As Secure-Instructsecure utilizes a much larger fine-tuning dataset (15.6K v.s. 856), it is not surprising that Secure-Instructsecure outperforms Secure-Instruct(vul, secure). To understand more about the comparison, we down‑sampled the Secure-Instructsecure dataset to 865 examples, while ensuring samples covering across all CWE-language, to match the size of the Secure-Instruct(vul, secure) dataset. Table 4 shows the comparison results across the four LLMs. Notably, the differences between the two are minimal, indicating essentially equivalent effectiveness. This shows that Secure-Instructsecure still generates highly effective synthetic data without being driven by vulnerable code first. It also indicates that there exist duplicates in the data by Secure-Instructsecure and opportunities for future improvement.
Summary of RQ1: Secure-Instructsecure significantly improves the studied LLMs in secure code generation, achieving an average increase of over 10%. When using tuning datasets of the same size, Secure-Instructsecure and Secure-Instruct(vul, secure) achieve comparable performance.
Model Metrics Pre-trained SafeCoder Secure-Instruct
(vul,secure) secure
CodeLlama- 7B Func@1 22.1% 14.9% 37.9% 34.9%
Sec@1 17.8% 17.4% 31.2% 33.7%
Func-Sec@1 8.9% 7.1% 22.9% 22.9%
CodeGen2- 7B Func@1 20.7% NA 20.0% 20.9%
Sec@1 17.9% NA 17.7% 16.4%
Func-Sec@1 7.4% NA 7.2% 7.5%
Mistral- 7B Func@1 30.1% 21.7% 34.4% 31.3%
Sec@1 20.4% 17.4% 25.5% 27.9%
Func-Sec@1 12.5% 11.5% 17.9% 18.3%
StarCoder- 1B Func@1 15.3% NA 13.8% 12.7%
Sec@1 16.2% NA 16.4% 15.5%
Func-Sec@1 6.1% NA 7.4% 7.4%
@@
Table 3: Func@1, Sec@1 and Func-Sec@1 results on CWEval for the pre-trained LLMs, SafeCoder fine-tuned LLMs, and Secure-Instruct (ours) fine-tuned LLMs. NA means Model not released by SafeCoder.
Models Secure-Instruct(vul, secure) Secure-Instructsecure
CodeLlama-7B 5,989/9,300 (64.4) 5,758/9,300 (61.9)
CodeGen2-7B 4,554/9,300 (49.0) 4,562/9,300 (49.1)
Mistral-7B 5,700/9,300 (61.3) 5,770/9,300 (62.0)
StarCoder-1B 5,900/9,300 (63.4) 6,157/9,300 (66.2)

### v2

@@ 3.3 Instruction Tuning LLMs for Secure Code Generation
Secure-Instruct utilizes LLMs to generate a problem description as an “instruction” for each of the instruction-tuning datasets. Then Secure-Instruct employs two instruction tuning methods, i.e., 1). Standard instruction tuning and 2). Secure instruction tuning.
@@
Generating Instruction. We follow the work (he2024instruction) to construct the instruction for each code snippet using LLMs. After that, we format the instruction-tuning datasets by putting the code snippet in the <<response>> tag and the generated instructions in the <<instruction>> tag, following a prior work (alpaca).
@@
Standard Instruction Tuning. This process fine-tunes an LLM on instruction-response pairs to improve its ability to follow natural language instructions. It uses supervised fine-tuning (SFT), where the model learns by minimizing loss over the provided dataset, which has been adopted by many prior studies (wei2024magicoder; xu2024wizardlm; luo2023wizardcoder).
@@
Secure-Specialized Instruction Tuning. SafeCoder (he2024instruction) proposed a specialized instruction-tuning technique for secure code generation by leveraging pairs of vulnerable code and fixed code mined from open-sourced software repositories. Since we collected 856 pairs of vulnerable and from Secure-Instruct(vul, secure), we applied this instruction tuning approach to our dataset. Prior research (he2023large) highlights a critical concern: exclusively using secure code examples for instruction-tuning can lead to a decline in code functionality. To address this limitation, we follow the work (he2024instruction) via the integration of a secure code dataset and a functional code dataset during instruction-tuning. It uses standard instruction tuning for the functional dataset and secure instruction tuning for the secure dataset to improve both functional and secure performance for LLMs. Specifically, we incorporate the functional code dataset from the Evol-Instruct dataset (luo2023wizardcoder), a high-quality dataset designed to enhance the functional robustness and generalization of code generation models.
4 Evaluation
4.1 Evaluation Setup
@@
Subject Large Language Models. We evaluate Secure-Instruct to improve four representative small and medium-sized LLMs with 1 to 7 billion parameters: CodeGen2-7B (nijkamp2023codegen2), CodeLlama-7B (roziere2023code), Mistral-7B (jiang2023mistral), and StarCoder-1B (li2023starcoder). These models are specifically designed for code generation, offering a balance between computational efficiency and performance. These models are widely used in prior works on secure code generation, such as (he2023large; he2024instruction; wei2024magicoder) for a fair comparison. For each of the four studied LLMs, we experimented with two settings of Secure-Instruct: (1) applying security-specialized instruction tuning on the dataset of pairs of vulnerable and secure code, i.e., from Secure-Instruct(vul, secure); and (2) applying standard instruction tuning on the secure code snippets from Secure-Instructsecure. Table 1 summarizes the statistics of the resulting fine-tuning datasets using Secure-Instruct.
Attribute Secure-Instruct Secure-Instruct
(vul, secure) secure
@@
Details on Instruction-Tuning LLMs. We use the parameter-efficient technique, LoRA (hu2021lora), to instruction-tune the 7B LLMs. LoRA reduces the GPU memory usage and maintains a promising performance. For the 1B model, we apply full-weight instruction-tuning. All experiments are conducted on five Nvidia RTX A6000 GPUs with 48GB of memory. For full-weight instruction-tuning, we set the learning rate to 2e-5, while for LoRA fine-tuning, the learning rate is 3e-4. Each model is tuned in 2 epochs. In LoRA, we configure the LoRA alpha to 16 to control the scaling of LoRA layers and set the dropout rate to 0.05.
Evaluation Benchmarks. We evaluate Secure-Instruct using two secure-code generation benchmarks: (1) CWEBench crafted and extended by us, and (2) CWEval (peng2025cweval), which enables validation of both functional correctness and security by test cases.
First, we craft CWEBench by extending prior benchmarks for secure code generation (siddiq2022securityeval; pearce2022asleep; he2024instruction). Such prior benchmarks are composed of security-specific scenarios and are sourced from three resources, i.e., MITRE documentation (CWE), CodeQL repository (codeql), and scenarios manually crafted by the authors. We first combined all three existing benchmarks (siddiq2022securityeval; pearce2022asleep; he2024instruction), removed duplicates, and excluded the scenarios derived from MITRE examples, as Secure-Instruct uses MITRE documents as seeds for LLMs to generate fine-tuning datasets. We define a secure ratio to measure LLM’s capability in CWEBench. Secure ratio is the percentage of secure code identified by CodeQL among all the generated code. To ensure robust results, we generated 100 samples for each security-specific scenario with a temperature of 0.8. A piece of LLM-generated code is deemed secure code if CodeQL detects no vulnerabilities in the code. Then, we followed a similar process as prior work (he2024instruction) to design 36 new security-specific scenarios covering 27 CWEs by adapting examples from the CodeQL repository. In the end, our crafted CWEBench consists of 93 security-specific scenarios covering 44 CWEs in six languages.
Second, we include a recent benchmark CWEval (peng2025cweval), which uses test cases for automated verification of both functional correctness and security without relying on static security checkers. In addition, CWEval provides manually crafted comprehensive specifications that describe complex user intentions and logic, which means it is a much closer step to real-world applications. In total, CWEval consists of 119 security scenarios covering 31 CWE types in five programming languages. CWEval comes with three metrics, namely Func@k, Sec@k and Func-Sec@k to evaluate the security and functionality of generated code. In this study, we set k as 1. Specifically, Func@1 measures the percentage of security scenarios whose top-1 generated code is functionally correct (i.e., passes all functional test cases). Sec@1 is defined as the percentage of security scenarios whose top-1 solution is secure (i.e., passes all security test cases). Func-Sec@1 concerns both functional correctness and security.
De-duplication between fine-tuning datasets and evaluation benchmarks. To prevent data leakage between the fine-tuning dataset and the two benchmark datasets, we performed de-duplication between training and evaluation following  (zhou2025lessleak). Specifically, we used MinHash with Locality-Sensitive Hashing (LSH) to detect potential duplicates. Code snippets were grouped based on similarity, and pairs with a Jaccard similarity above 0.7 were flagged as potential duplicates. After that, we removed these flagged snippets from the synthesized dataset.
4.2 Evaluation Results
In this section, we present our evaluation results in both secure and functional benchmarks.
@@ 4.2 Evaluation Results
Results of CWEBench. We evaluated the pre-trained models and instruction-tuned models by Secure-Instruct. We also conducted a comparative analysis between Secure-Instruct and SafeCoder (he2024instruction). In particular, we reused the two released models, i.e., CodeLlma-7B and Mistral-7B, which were already instruction-tuned by SafeCoder. Table 2 presents the results on CWEBench. The highest secure ratio is highlighted in bold. Across all models, Secure-Instruct(vul, secure) and Secure-Instructsecure demonstrate significant improvements in secure code generation. Compared with the pre-trained models, Secure-Instruct achieved the highest secure ratio across all models. For example, CodeLlama‑7B achieved the highest secure ratio among the four models when instruction‑tuned with Secure-Instructsecure, improving by 46.6% (the secure ratio from 47.6% to 69.8%), compared to an increase 35.3% under Secure-Instruct(vul, secure). The secure ratio of CodeGen2-7B is improved from 48.8% to 57.0% of secure ratio by Secure-Instructsecure and to 49.0% by Secure-Instruct(vul, secure). Mistral-7B achieved an increase of 28.9% by Secure-Instructsecure. StarCoder-1B also achieves a notable improvement, i.e., secure ratio rising from 55.4% to 67.5%. Across the four LLMs, Secure-Instruct improves them to generate more secure code, and Secure-Instructsecure shows consistently better results than Secure-Instruct(vul, secure).
Compared to SafeCoder, our two models consistently outperform the SafeCoder models, with +15.2% in CodeLlama-7B and +9.9% in Mistral-7B using Secure-Instructsecure. At the CWE category level, CodeLlama‑7B instruction tuned with Secure-Instructsecure achieves the highest or ties for highest secure ratio in 29 of 44 CWEs, outperforming both its pre-trained baseline and the SafeCoder-tuned variant. Similarly, Mistral‑7B with Secure-Instructsecure leads in 25 of 44 CWEs relative to its pretrained and SafeCoder‑tuned counterpart. Due to the limited space, we put all detailed analyses in our replication package. These results demonstrate that Secure-Instructsecure significantly enhances secure code generation, outperforming the SOTA approach, SafeCoder.
Code Llama- 7B Code Gen2- 7B Mistral- 7B Star Coder- 1B
Pre-trained 47.6% 48.8% 51.5% 55.4%
SafeCoder 60.6% ($\uparrow$ 27.3%) NA 60.4% ($\uparrow$ 17.3%) NA
Secure-Instruct (vul,secure) 64.4% ($\uparrow$ 35.3%) 49.0% ($\uparrow$ 0.4%) 61.3% ($\uparrow$ 19.0%) 63.4% ($\uparrow$ 14.4%)
Secure-Instruct secure 69.8% ($\uparrow$ 46.6%) 57.0% ($\uparrow$ 16.8%) 66.4% ($\uparrow$ 28.9%) 67.5% ($\uparrow$ 21.8%)
Table 2: The secure ratio for pre-trained and instruction-tuned models using Secure-Instruct and SafeCoder on CWEBench. $\uparrow$ and $\downarrow$ indicate relatively change from pre-trained models. NA means the model not released by SafeCoder.
@@
Results of CWEval. Table 3 presents the results on the benchmark CWEval. Similar to CWEBench, Secure-Instruct achieves the highest Func-Sec@1 across all evaluated LLMs. Notably, CodeLlama-7B and Mistral-7B show significant improvements, with increases of 157.3% and 46.4%, compared to their pre-trained versions. For Sec@1, CodeLlama-7B, CodeGen2-7B, and StarCoder-1B achieved the highest secure ratio using Secure-Instruct, with 33.7%, 27.9%, and 16.4%, respectively. The results of Func@1 are put in our replication package due to limited space.
Interestingly, SafeCoder exhibits a performance drop on CWEval. For example, SafeCoder’s CodeLlama-7B achieves only 7.1% Func-Sec@1, which is lower than the pre-trained model. A similar decline is observed in SafeCoder’s Mistral-7B. These findings are consistent with the observations reported in the study (peng2025cweval). Overall, LLMs with Secure-Instruct consistently outperform both their pre-trained versions and SafeCoder models in secure code generation across CWEBench and CWEval.
CodeLlama-7B CodeGen2-7B Mistral-7B StarCoder-1B
Sec@1 Func-Sec@1 Sec@1 Func-Sec@1 Sec@1 Func-Sec@1 Sec@1 Func-Sec@1
Pre-trained 17.8% 8.9% 17.9% 7.4% 20.4% 12.5% 16.2% 6.1%
SafeCoder 17.4% ($\downarrow$ 2.2%) 7.1% ($\downarrow$ 20.2%) NA NA 17.4% ($\downarrow$ 14.7%) 11.5% ($\downarrow$ 8.0%) NA NA
Secure-Instruct (vul,secure) 31.2% ($\uparrow$ 75.3%) 22.9% ($\uparrow$ 157.3%) 17.7% ($\downarrow$ 1.1%) 7.2% ($\downarrow$ 2.7%) 25.5% ($\uparrow$ 25.0%) 17.9% ($\uparrow$ 43.2%) 16.4% ($\uparrow$ 1.2%) 7.4% ($\uparrow$ 21.3%)
Secure-Instruct secure 33.7% ($\uparrow$ 89.3%) 22.9% ($\uparrow$ 157.3%) 16.4% ($\downarrow$ 8.4%) 7.5% ($\uparrow$ 1.4%) 27.9% ($\uparrow$ 36.8%) 18.3% ($\uparrow$ 46.4%) 15.5% ($\downarrow$ 4.3%) 7.4% ($\uparrow$ 21.3%)
@@
Table 3: Sec@1 and Func-Sec@1 results on CWEval for pre-trained LLMs and LLMs tuned by SafeCoder and Secure-Instruct (ours).
Results of Functional Benchmarks. We evaluated our approaches with functional benchmarks, i.e., HumanEval and MBPP, two widely used benchmarks for assessing code generation in Python. For each task, we generate 40 samples and assess LLMs’ functionality using the Pass@k metric, including Pass@1. Pass@k is a standard metric for evaluating code generation, which represents the probability that at least one of the $k$ generated candidates for a given problem is correct. Table 4 presents the Pass@1 (Pass@10 is available in our replication package) on the HumanEval and MBPP benchmarks for pre-trained LLMs, instruction-tuned LLMs by SafeCoder and Secure-Instruct. For CodeLlama‑7B, using Secure-Instructsecure achieves the highest Pass@1 on HumanEval at 35.6%, outperforming both the pre-trained and SafeCoder versions. For Mistral‑7B, the model using the Secure-Instruct(vul, secure) achieves the highest score on MBPP at 30.7%. Codegen2-7B achieved the highest Pass@1 using Secure-Instructsecure, with 16.3% on HumanEval and 22.8% on MBPP. Overall, most models show performance improvements on both datasets, except for StarCoder-1B.
CodeLlama-7B CodeGen2-7B Mistral-7B StarCoder-1B
HumanEval MBPP HumanEval MBPP HumanEval MBPP HumanEval MBPP
Pre-trained 25.5% 28.3% 15.3% 20.4% 23.0% 24.7% 11.7% 13.7%
SafeCoder 32.0% ($\uparrow$ 25.5%) 33.7% ($\uparrow$ 19.1%) NA NA 31.3% ($\uparrow$ 36.1%) 29.3% ($\uparrow$ 18.6%) NA NA
Secure-Instruct (vul,secure) 34.7% ($\uparrow$ 36.1%) 36.0% ($\uparrow$ 27.2%) 8.1% ($\downarrow$ 47.1%) 15.4% ($\downarrow$ 24.5%) 31.6% ($\uparrow$ 37.4%) 30.7% ($\uparrow$ 24.3%) 6.7% ($\downarrow$ 42.7%) 12.2% ($\downarrow$ 10.9%)
Secure-Instruct secure 35.6% ($\uparrow$ 39.6%) 34.7% ($\uparrow$ 22.6%) 16.3% ($\uparrow$ 6.5%) 22.8% ($\uparrow$ 11.8%) 25.9% ($\uparrow$ 12.6%) 27.1% ($\uparrow$ 9.7%) 10.9% ($\downarrow$ 6.8%) 17.7% ($\uparrow$ 29.2%)
Table 4: Pass@1 results on HumanEval and MBPP for pre-trained LLMs and LLMs tuned by SafeCoder-tuned and Secure-Instruct.
5 Discussions
The accuracy of CodeQL. In CWEBench, the ground-truth of security is based on CodeQL, which is a static tool and may produce false positives. To address this concern, we randomly sampled 291 code snippets labeled as secure by CodeQL from four models and manually verified their security. All sampled snippets were confirmed to be truly secure. To further support our conclusion, we used an additional benchmark, CWEval, which evaluates security through test cases.
The similarity between Secure-Instruct’s instruction-tuning datasets and our evaluation benchmark. We performed a similarity analysis between the instruction-tuning data and benchmark data. Specifically, we first extracted all the secure code examples from the CodeQL repository for the 44 CWEs of our benchmark. We compare each LLM-generated secure code with the secure code examples from the CodeQL repository under the same CWE. We then calculated the cosine similarity of each pair using TF-IDF embeddings (sparck1972statistical). Figure 5 depicts the average cosine similarity score for the data generated by Secure-Instructsecure, which is 0.17, while the average score for data generated by Secure-Instruct(vul, secure) is 0.23. Secure-Instructsecure’s values align with findings from previous work (wei2024magicoder) on synthesizing instruction-tuning data using LLMs, which reported similarity scores typically ranging between 0.1 and 0.2 in the previous synthesis techniques. Importantly, this result indicates that the improvements from Secure-Instruct are not simply due to the inclusion of data from the same distribution of the CodeQL repository. Instead, the generated data provides additional diversity for the instruction-tuning dataset.
Figure 4: Cosine similarities between Secure-Instruct’s instruction-tuning dataset and the secure code examples by CodeQL (i.e., our evaluation benchmark).
Low cost of Secure-Instruct in generating instruction-tuned datasets. Secure-Instruct generates high-quality instruction-tuning data with low cost and no human effort. Using ChatGPT (GPT-4o-2024-08-06), the total cost for generating all secure code snippets is USD 114.09. Similarly, the cost is USD 109.74 for Claude. The cost per CWE-language is USD 2.87. In contrast, SafeCoder (he2024instruction) faces scalability challenges, filtering 145M commits down to just 1,200 code pairs, and requiring manual selection of 465 high-quality examples. Secure-Instruct eliminates such overhead, offering a faster, cheaper, and automated alternative for generating secure and reliable datasets.
Code Llama- 7B Code Gen2- 7B Mistral- 7B Star Coder- 1B
Secure-Instruct(vul, secure) 64.4% 49.0% 61.3% 63.4%
Secure-Instructsecure 61.9% 49.1% 62.0% 66.2%